OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (-5, 1) with a slope of 3/2. A) y - 5 = 3/2(x + 1) B) y + 1 = 3/2(x - 5) C) y + 5 = -3/2(x - 1) D) y - 1 = 3/2(x + 5)

OpenStudy (ghazi):

y-1= 3/2*(x-(-5))

OpenStudy (ghazi):

it seems none of your options are correct

OpenStudy (karatechopper):

Lemme try to solve?

OpenStudy (anonymous):

really? these are the one's i have listed though..? @ghazi

OpenStudy (karatechopper):

.....Ghazi is right

OpenStudy (ghazi):

you just need to use standard equation of line (y-y1)=slope *(x-x1) and on the basis of this i can't see any of the option matching the answer

OpenStudy (anonymous):

okay how about now guys?! @karatechopper @ghazi i think my computer deleted the fractions

OpenStudy (ghazi):

it's obviously D

OpenStudy (karatechopper):

Hahaha! Now it makes sense! Ok its the last one. Thats the one i got the answer for